In another 3 years Tommy will be 3 times as old as he was 3 years ago. How old is he now?
Answer by ptfile(81)

Let x+3 = Tommy's age in 3 years
x = Tommy's current age
x-3 = Tommy's age 3 years ago
In another 3 years, Tommy will be 3 times as old as he was 3 years ago.
x+3 = 3(x-3)
x+3 = 3x-9
x = 3x-12
12 = 2x
6 = x
Check
6+3 = 3(6-3)
9 = 3(3)
9 = 9